Laws of exponents to simplify
not sure if I will write this out correctly but I'll give it a whirl!
I also worked it and will provide my answer! Is there a way to check your answer?
question? Ohh, the # is powered^ by a fraction.
3^3/4 * 3^7/8
my solution:
3^3/4 * 3^7/8= 3 3/4*7/8= 3 36/32, but I feel that I need to simplify further? break the fraction of 36/32 down.
Am I getting the process down? I don't know why it is so confusing!
Thanks Again!
Julie :)

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

3^3/4 * 3^7/8
------------
Same base so add the exponents.
------------
= 3^(3/4 + 7/8)
---
Find a common denominator.
= 3^(6/8 + 7/8)
---
Add the exponents.
= 3^(13/8)
------------
Cheers,
Stan H.
===========

Answer by solver91311(24713)



Fortunately for you right now and for all of us at one time or another, violating the Laws of Exponents is not a criminal offense. If it were, I'm afraid your penance would be long and severe.



You multiplied where you should have added.

Perhaps you were thinking of this situation:



When raising a power to a power -- then you multiply the exponents.

For your problem:




John